OpenStudy (anonymous):

Which mathematical sentence represents the solution dor d in the equation 6e = ef + 3d A. d = 6e - ef/3 B. d = 6e + ef/3 C. d = 2e - ef/3 D. d = 2e + ef/3 Helpp ! i know its not D

OpenStudy (anonymous):

6e = ef + 3d subtract ef from each side

OpenStudy (anonymous):

Can you explain a little bit more ? Im still confused

OpenStudy (anonymous):

6e = ef + 3d - ef - ef 6e - ef = ef -ef +3d 6e - ef = 3d divide by 3 2e - ef/3 =d

OpenStudy (anonymous):

hmm okay, thank you